OpenStudy (anonymous):

Determine the point on the hyperbola −7x2+6y2=20closest to the point (-2, 0).

OpenStudy (anonymous):

let (h,k) be the points. so the equation of tangent on hyper bola at this point -7hx+6ky=20 now slope of tangent= 7h/6k. and slope of line joining (-2,0) and (h,k)..=(0-k)/(-2-h) both line are perpendicular as the distence is shortest.so. (7h/6k).(-k/-2-h)=-1. h=-12/13. put the value in equation as both h k lies on hyperbola... -7(-12/13)^2+6k^2=20 from here you can solve the value of k